OpenStudy (anonymous):

how do i solve for x in the following, 7x-5+2x=3-4x+12

OpenStudy (anonymous):

move all the x terms to one side .. to the left side for example you do this by adding 4x to both sides of the equation what will you get ?

OpenStudy (anonymous):

9x-5=-4x+15?

OpenStudy (anonymous):

no 7x-5+2x=3-4x+12 adding 4x to both sides -> 7x +2x+4x -5 = 3+12-4x+4x 13x-5=15

OpenStudy (anonymous):

all the terms of x. 9x+4x=20

OpenStudy (anonymous):

13x=20 x=20/13

OpenStudy (anonymous):

nok thank you. i had that all messed up

OpenStudy (anonymous):

now we need to move this '5' to the right side since we want all the x terms at the left sides .. we do this by adding 5 to both sides of the equation what we get?

OpenStudy (anonymous):

@miteshchvm we better let people try and learn then solve for them ..

OpenStudy (anonymous):

than*

OpenStudy (anonymous):

ya dude @Coolsector ;) sure

OpenStudy (jiteshmeghwal9):

take the variables at one side & constants at one side as\[7x+2x+4x=3+12+5 .........(By \space Transposition )\]\[13x=20\]by dividing 13 both sides.we have\[\frac{13x}{13}=\frac{20}{13} \implies x=\frac{20}{13}\]

OpenStudy (jiteshmeghwal9):

gt it @Jesusfelon ?